OpenStudy (chelsie):

For the function f(x)=f(x)=(7-8x)^2, find f^-1. Determine whether f^-1 is a function.

OpenStudy (prathamesh_m):

To find the inverse of a function, consider: \[y = (7-8x)^{2}\] Now find x as a function of y from the above equation.(as in x = f(y))

OpenStudy (prathamesh_m):

Let the equation you get once you do that be x = g(y). Then g(y) is the inverse, i.e., \[g(y) = f ^{-1} (y)\] To check if its a function, plot its graph and check whether at any point there are two values of y for a single value of x.

OpenStudy (zzr0ck3r):

Here is an example of something similar. \(f(x) = x^2-1\) First make it \(y=x^2-1\) and now solve in terms of \(x\) \(y+1=x^2\implies x=\pm\sqrt{y+1}\) Surely this is not a function as we get two answers for every \(y>-1\)